OpenStudy (mathhelp346):

Does -1/2m(4m^2-2mn)=-2m^3+m^2n?

OpenStudy (lgbasallote):

is that \[\huge -\frac 1{2m} (4m^2 - 2mn)= -2m^3 + m^2n\]

OpenStudy (mathhelp346):

yes except the m isnt in the denominator

OpenStudy (lgbasallote):

ooh...well then \[\huge -\frac 12 m ( 4m^2 - 2mn)\] distribute \[\huge \implies -\frac{4m^2 \times m}2 - \frac{-2mn \times m}2\] multiply.. \[\huge \implies -2m^3 + m^2 n\] so yeah you're right

OpenStudy (mathhelp346):

ok thanks

OpenStudy (lgbasallote):

welcome

OpenStudy (mathhelp346):

lol thanks